Wendy was awarded a volleyball scholarship to the University of Michigan, so on graduation her parents gave her the $22,000 they had saved for her college tuition. She opted to invest some money in a privately held company that pays 10 % per year and evenly split the remaining money between a money market account yielding 2 % and a high-risk stock that yielded 40 %. At the end of the first year she had $26,400 total. How much did she invest in each of the three?
Answer by jorel1380(3719)

Let n be the amount of money invested at 10% per year. Then the remaining money is 22000-n. This gets split into two; being .5(22000-n) and another .5(22000-n). If she has $26400 after her first year, she has made 26400-22000=$4400 interest. So:
.1(n)+.02(.5(22000-n)+.4(.5(22000-n)=4400
4620-.11n=4400
.11n=220
n=2000
22000-n=20000
20000/2=10000
$2000 was invested at 10%; and $10000 was invested at 2% and 40% each. ☺☺☺☺